# MuSACo: Multimodal Subject-Specific Selection and Adaptation for Expression Recognition with Co-Training

### v1

@@ 6.6 MuSACo
MuSACo (inspired by the layered harmony of Musaca) integrates two synergistic modules for subject-specific adaptation: a co-training-based source subject selection module that identifies the most relevant sources using complementary cues from multiple modalities, and an adaptation module that aligns source and target domains using class-aware and class-agnostic losses. Like Musaca’s layered composition—where distinct elements work in harmony—MuSACo leverages modality-specific strengths to guide pseudo-label generation and fuse information effectively. This enables robust, personalized adaptation for each target subject while maintaining consistency across modalities.

### v2

@@ 13 MuSACo
MuSACo (inspired by the layered harmony of Musaca) integrates two synergistic modules for subject-specific adaptation: a co-training-based source subject selection module that identifies the most relevant sources using complementary cues from multiple modalities, and an adaptation module that aligns source and target domains using class-aware and class-agnostic losses. Like Musaca’s layered composition—where distinct elements work in harmony, MuSACo leverages modality-specific strengths to guide pseudo-label generation and fuse information effectively. This enables robust, personalized adaptation for each target subject while maintaining consistency across modalities.